Find the rational numbers between -2/5 and ½. 
Find the rational numbers between -2/5 and ½. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us make the denominators same, say 50.
-2/5 = (-2 × 10)/(5 × 10) = -20/50
1⁄2=(1×25)/(2×25)=25/50
Ten rational numbers between -2/5 and 1⁄2 = ten rational numbers between -20/50 and 25/50
Therefore, ten rational numbers between -20/50 and 25/50 = -18/50, -15/50, -5/50, -2/50, 4/50, 5/50, 8/50, 12/50, 15/50, 20/50

Thanks